OpenStudy (anonymous):

If the 2 lines given by the equations 12x+3y=20 and ax+y=10 don't intersect when graphed, what is the value of "a"?

OpenStudy (sasogeek):

if they don't intersect, it means they're parallel and their gradient is the same. find the gradient of the first equation and that's equal to "a" :)

OpenStudy (zzr0ck3r):

set both equal to y, and make them have the same slope

OpenStudy (zzr0ck3r):

as long as they dont have the same y intercept they will not cross if they have the same slope

OpenStudy (zzr0ck3r):

3y = 20 -3x y = -x-(20/3) and the other equation set to y is what?

OpenStudy (zzr0ck3r):

@mitchelsewbaran ?

OpenStudy (anonymous):

hmmm.....

OpenStudy (zzr0ck3r):

that should be +(20/3)

OpenStudy (zzr0ck3r):

y = -x+(20/3)

OpenStudy (zzr0ck3r):

solve for y in the second equation

OpenStudy (zzr0ck3r):

y = -ax+10 we want them to have the same slope so what does a need to be? y = -x+(20/3) y = -ax+10 a = ?

OpenStudy (anonymous):

10?

OpenStudy (sasogeek):

12x+3y=20 solve for y 3y=-12x+20 y=-4x+(20/3) can u tell what the slope is?

OpenStudy (sasogeek):

@mitchelsewbaran ... ?

OpenStudy (anonymous):

hold on